Problem (10 points: Nike, Inc. is considering a $10M investment paid in smarter Sneakers The new sneakers are expected to ge need show work process not excel.
Answer #1

Calculation of W ACC for profect & Cost of Equity = 8 + (80 %) x Beta = % a + 5X 1.2 2x (1-0.4) = = 8 1.2 - % WACC = ke xwe +